Using the method of telescopic sums, solve 

the recurrence relation : (n+1)xn-nxn-1-n/2,x0=10,n≥1


1
Expert's answer
2021-08-03T07:06:17-0400

"(n+1)x_n-nx_{n-1}=n\/2"

Let "nx_{n-1}=y_n". Then "y_{n+1}-y_n=n\/2" and "y_1=x_0=10".

"y_n=(y_n-y_{n-1})+(y_{n-1}-y_{n-2})+\dots+(y_2-y_1)+y_1="

"=\frac{n-1}{2}+\frac{n-2}{2}++\dots+\frac{1}{2}+10=10+\frac{n(n-1)}{4}"

"x_{n-1}=y_n/n=\frac{10}{n}+\frac{n-1}{4}"

"x_n=\frac{10}{n+1}+\frac{n}{4}"

Check the solution:

"x_0=\frac{10}{0+1}+\frac{0}{4}=10"

"(n+1)x_n-nx_{n-1}=(n+1)(\frac{10}{n+1}+\frac{n}{4})-n(\frac{10}{n}+\frac{n-1}{4})="

"10+\frac{n(n+1)}{4}-(10+\frac{n(n-1)}{4})=n/2"